# VLM-Social-Nav: Socially Aware Robot Navigation through Scoring using Vision-Language Models

### v2

@@ I Introduction
Mobile robots integrated into diverse indoor and outdoor human-centric environments are becoming increasingly prevalent. These robots serve various functions, ranging from package and food delivery [starship, scout] to service [dilligent] and home assistance [astro]. Overall, these roles necessitate interaction with humans and navigating seamlessly through public spaces with pedestrians. In such dynamic scenarios, it is important for the robots to engage in socially compliant interactions and navigation [xiao2022motion, mirsky2024conflict, mavrogiannis2023core].
@@
Inferring contextually appropriate navigation behaviors is challenging. Humans have various behaviors and the environmental or task contexts also cannot be easily categorized [mavrogiannis2023core]. A common strategy to handle the challenge is by learning-based approaches to empirically learn the complicated contexts. Imitation Learning (IL) is a recent emerging paradigm for desired navigation behavior [hirose2023sacson, xiao2020appld, xiao2022learning, panigrahi2023study, raj2024targeted]. This approach enables autonomous robots to navigate socially by learning from human demonstrations. Other learning approaches, such as Reinforcement Learning (RL) have also been used to address this problem [kretzschmar2016socially]. While both methods demonstrate promising results in real-world settings, substantial datasets [thorDataset2019, karnan2022scand, nguyenmusohu] for training and reward engineering are required for their successful application and it is hard to generalize.
Language models are inherently well-suited for contextual understanding but not well applied in social navigation. Recent Large Language Models (LLMs) and Vision-Language Models (VLMs) demonstrate a deep understanding of contextual information and have the potential to perform chain-of-thought [wei2022chain] and common sense reasoning [geva2021did, gpt4v2023]. Those processes are inherent to social navigation, especially the challenges of contextual appropriateness and politeness, which require understanding the task/environmental context and the behavior of humans. This capability has also been evaluated across diverse domains of robotics, including human-like driving scenarios [wen2024road, sha2023languagempc] and autonomous robot navigation [10160969, shah2022lmnav]. However, using language models for social navigation is not well explored and the language models suffer from high latency for real-time navigation, and the issue impedes the smoothness and efficiency of human-robot social interaction.
@@
Figure 1: The trajectories of our approach, VLM-Social-Nav (red), DWA (blue), and BC (yellow) approaches in the frontal encountering scenario (left) and the intersection scenario (right). The resulting trajectories show that VLM-Social-Nav demonstrates more social compliant behavior as it is instructed by a prompt (i.e., Move to the right when passing by a person, and do not obstruct others’ paths).
@@
Main Results: In this paper, we present VLM-Social-Nav, a new approach that uses VLMs to interpret contextual information from robot observation to help autonomous robots improve their navigation abilities in human-centered environments. In particular, we leverage a VLM to analyze and reason about the current social interaction, and generate an immediate preferred robot action to guide an underlying motion planner. Our VLM-based scoring module computes a social cost, which is used for the bottom-level motion planner to output the appropriate robot action. To overcome the limitation of existing VLMs’ latency issue, we utilize a state-of-the-art perception model (i.e., YOLO [YOLO2023]) to detect key entities that are used for social interactions (e.g., humans, gestures, and doors) and query a VLM to generate socially compliant navigation behavior and compute the social cost. We demonstrate VLM-Social-Nav in four different indoor scenarios with human interactions. Unlike previous social navigation approaches, VLM-Social-Nav can better navigate through social scenarios by interpreting the situation based on common sense without any dedicated training on a large dataset. Some of our main results include:
@@
  • We propose VLM-Social-Nav, a novel approach for social robot navigation by integrating VLMs with optimization-based or scoring-based motion planners and a state-of-the-art perception model for better VLM efficiency.
@@
  • We evaluate VLM-Social-Nav in four different real-world indoor social navigation scenarios along with a user study and compare the results with a Dynamic-Window Approach (DWA) [fox1997dynamic] and Behavior Cloning (BC) [pomerleau1988alvinn] method trained on a state-of-the-art large Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand]. VLM-Social-Nav achieves at least $27.38\%$ improvement in average success rate and $19.05\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@ II-A Safety Requirement of Social Navigation
For social navigation, safety is a basic requirement for interaction with humans and navigation in dynamic challenging scenarios [van2010optimal, fox1997dynamic, liang2021vo, liang2021crowd]. DWA [fox1997dynamic] calculates the collision constraints of the robots’ actions and chooses the best feasible action closest to the target as the output. Model Predictive Control (MPC) approaches [10341869, williams2017information] are also widely used for collision avoidance tasks, which provide smooth trajectories for the robot to follow; learning-based MPC [williams2017information, xiao2022learning] enhances the performance of the original MPC method by empirically learning the best actions for navigation, which requires a large expert demonstration dataset for training. Velocity-Obstacle (VO)-based approaches are more efficient and can be used to simulate the actions of crowds [van2010optimal, best2014densesense], but these approaches don’t take into account the uncertainties in the perception output. PRVO [gopalakrishnan2017prvo] and OFVO [liang2021vo] handle the uncertainties of perception in motion planning, but those approaches require a hard threshold to set the confidence for planning. To deal with this issue, learning-based methods empirically train the policies by demonstrations [liang2021crowd, aradi2020survey, sun2021motion], whereas other methods [liang2021crowd, aradi2020survey] use reinforcement learning to train the robot in a simulator and implement it in real-world scenarios. However, learning-based approaches require a significant amount of data or on-policy training with realistic simulators to learn the task. Social interactions are highly nuanced and context-dependent. Simulating these interactions accurately requires sophisticated models of human behavior and interaction dynamics, which is not trivial.
@@ II-B Contextual Appropriateness of Social Navigation
Researchers have proposed diverse methodologies to integrate social awareness into mobile robot navigation systems. The development of a comprehensive social navigation system poses inherent complexity, requiring sophisticated perception and reasoning capabilities to navigate environments shared with humans and other robots [mirsky2024conflict]. Establishing an accurate definition of social navigation is paramount, given its potential variation across cultures and platforms. Except for the safety requirement, assessing social compliance presents an additional challenge, contingent on the scenario and platform, and should consider contextual appropriateness [gao2022evaluation, francis2023principles]. Various methodologies are employed to address this challenge, with a significant focus on enhancing learning methods through reinforcement learning, learning from demonstration—particularly by analyzing examples of human trajectories or robots operated by humans—and the utilization of simulated datasets [9679193, 5354147, 7539621, li2017role, nazeri2024vanp]. Additionally, various datasets have been collected for this purpose. Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand] and Multi-Modal Social Human Navigation Dataset (MuSoHu) [nguyenmusohu] are two recent large-scale social human navigation datasets in many natural human-inhabited public spaces for robots to learn similar, human-like, socially compliant navigation behaviors. Despite extensive research employing diverse approaches with machine learning techniques [10323465], VLMs have not been used to solve the social navigation problem, which, however, have shown the potential to successfully analyze the contextual information of the environment.
@@ II-C Large Foundation Models for Navigation
Recent advancements in Language Foundation Models (LFMs) [bommasani2022opportunities], encompassing Vision-Language Models (VLMs) and Large-Language Models (LLMs), show significant potential in robotic navigation, despite the challenges like bias and reliability [payandeh2023susceptible] associated with the LFMs. SayCan [Ahn2022] integrates LLMs for high-level task planning. GPT-Driver [mao2023gptdriver] evaluates the performance of GPT-3.5 in simulation for autonomous driving, framing motion planning as a language modeling problem. L3MVN [yu2023l3mvn] constructs semantic maps of environments and utilizes LLMs to reach long-term goals, while LLaDA [li2024driving] enables autonomous vehicles to adapt to diverse traffic rules across regions. LM-Nav [shah2022lmnav] utilizes GPT-3 and CLIP [radford2021learning] to navigate outdoor environments based on natural language instructions, combining language and visual cues for optimal path planning. Despite their powerful capabilities in contextual understanding and commonsense reasoning, language models have not been extensively investigated for social navigation. Our approach proposes a novel method to navigate robots in a socially compliant manner.
@@ III-A Problem Definition
For social robot navigation, humans are no longer perceived only as dynamic obstacles but also as social entities [rios2015proxemics, mirsky2024conflict]. It necessitates integrating social norms into robot behaviors. We define the social robot navigation problem as a Markov Decision Process (MDP): $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C}\rangle$ where $\mathbf{s}=(x,y,\theta)\in\mathcal{S}$ is a state consisting of a robot pose, $\mathbf{a}=(v,w)\in\mathcal{A}$ is an action consisting of a linear and an angular velocity of the robot, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function characterizing the dynamics of the robot, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a cost function. Given a cost function $\mathcal{C}$, the motion planner finds $(v^{*},w^{*})$ that minimizes the expected cost. The cost function takes the following form:
@@
where $\mathcal{C}_{\textrm{goal}}$ encourages movement toward the goal, $\mathcal{C}_{\textrm{obst}}$ discourages collisions with obstacles, and $\mathcal{C}_{\textrm{social}}$ encourages the robot to follow the social norms. $\alpha$, $\beta$, and $\gamma$ are a non-negative weight for each cost term.
@@
where $\mathcal{B}$ is a navigation behavior, and $\mathcal{B}_{h}$ is a navigation behavior humans would adopt in accordance with social conventions. Minimizing the deviation between them will encourage the robot to emulate socially acceptable human behaviors. While $\mathcal{B}_{h}$ can be obtained through various methods, including large datasets [thorDataset2019, karnan2022scand, nguyenmusohu], we leverage the power of a VLM to compute appropriate behavior based on its rich contextual understanding and nuanced interpretations from perceived images and given prompts. We elaborate further in Section III-C.
@@ III-B VLM-based Social Navigation Architecture
Figure 2 highlights the overview of VLM-Social-Nav. Our formulation is built upon an autonomous navigation system comprising a perception layer and an optimization-based motion planner. The motion planner considers the sensor inputs and outputs a robot action that minimizes the cost function $C$.
While LiDAR detects geometric information useful for obstacle avoidance, RGB images provide contextual details of the current environment. They contain rich information crucial for social navigation. To enhance navigation capabilities within social contexts, we propose a VLM-based scoring module. VLMs excel in contextual understanding, interpreting scenes not solely based on visual features but also considering social dynamics. VLMs generate socially appropriate robot actions based on current observations and input instructions. Our VLM-based scoring module then calculates a cost term to be used by the motion planner.
@@ 5       $e\leftarrow\text{Perception Model}(\mathcal{I})$;
6       $(\delta_{s},\delta_{d})\leftarrow 0$;
@@ 8             $\mathcal{B}_{h}=\text{VLM}(\mathcal{I},\mathcal{P},\mathbf{a})$;
9             $\mathcal{B}_{h}\mapsto(\delta_{s},\delta_{d})$;
10
11       forall possible actions $\mathbf{a}$ do
12             $\mathcal{C}_{\textrm{social}}=\text{VLM-based scoring}(\delta_{s},\delta_{d})$;
13             Calculate the total cost $\mathcal{C}=\alpha\cdot\mathcal{C}_{\textrm{goal}}+\beta\cdot\mathcal{C}_{\textrm{obst}}+\gamma\cdot\mathcal{C}_{\textrm{social}}$;
14
@@ 17       Update robot position $\mathbf{p}_{r}$;
Algorithm 1 VLM-Social-Nav
@@
Figure 3: An example input image ($\mathcal{I}^{t}$) and prompt ($\mathcal{P}$) used in VLM-Social-Nav. Parameterized inputs ($\mathbf{a}^{t}$) are highlighted in blue. Formatted output specifying the heading ($\delta_{d}$) and the speed ($\delta_{s}$) are highlighted in red. The example input data is one of the frontal approach scenarios from MuSoHu [nguyenmusohu]. The output of VLM is Move right with slow down, which was the same action the collected data took.
III-C VLM-based Scoring Module
VLM plays a crucial role in VLM-Social-Nav to infer immediate socially compatible navigation behavior $\mathcal{B}_{h}^{t+1}$ based on its pre-trained large internet-scale dataset:
where $\mathcal{I}^{t}$ is an RGB image from the robot view at time $t$, $\mathcal{P}$ is a textual prompt, $\mathbf{a}^{t}$ is a current robot action at time $t$. Inspired by In-Context Learning (ICL), our prompt $\mathcal{P}$ is designed to leverage the VLM’s reasoning abilities through zero-shot examples. This approach offers an interpretable interface, mirroring human reasoning and decision-making processes, without extensive training [min2022rethinking].
where $w_{l}$ and $w_{a}$ are non-negative weight values. Given all the cost terms, our low-level optimization-based motion planner finds the robot action $(v^{*},w^{*})$ that minimizes the cost.
Figure 3 shows an example prompt $\mathcal{P}$ used in our experiment. We provide a high-level task description along with an image $\mathcal{I}^{t}$ captured from the robot’s perspective. Furthermore, the current robot action $\mathbf{a}^{t}=(v^{t},w^{t})\in\mathcal{A}$ is provided. Supplementary instructions regarding walking etiquette are also included. Although the VLM demonstrates proficient navigation abilities even in the absence of explicit instructions, offering reasoning guidelines enhances its decision-making processes [min2022rethinking]. These guidelines not only facilitate comprehensive reasoning and judgment within the VLM but also enable the robot to adapt to specific rules more effectively. For example, in a country where it’s customary to walk on the left, we can rephrase the prompt as “Move to the left when passing by another person”.
@@ III-C VLM-based Scoring Module
Figure 4: Qualitative Results: the robot navgation behaviors with VLM-Social-Nav for four social navigation scenarios: (a) Frontal Approach, (b) Frontal Approach with Gesture, (c) Intersection, and (d) Narrow Doorway. The gray solid arrow shows the participant’s path. The red solid arrow shows the robot’s path. The red and gray dashed arrows show the robot’s and participant’s path respectively after a stop motion. A caption on the top left shows the result from the VLM.
TABLE I: Quantitative Results: performance comparisons using BC [pomerleau1988alvinn] trained on SCAND dataset [karnan2022scand], DWA [fox1997dynamic], and VLM-Social-Nav. VLM-Social-Nav achieves at least $27.38\%$ improvement in average success rate and $19.05\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
Metric Method Scenario
Success Rate (%) BC 38.10 0 33.33 42.86
DWA 100 0 90.48 100
VLM-Social-Nav 100 100 100 100
Collision Rate (%) BC 42.86 66.67 28.57 38.10
DWA 28.57 19.05 19.05 38.10
VLM-Social-Nav 14.29 0 4.76 9.52
User Study Score BC 2.80 $\pm$ 1.45 2.23 $\pm$ 1.54 2.80 $\pm$ 1.40 2.60 $\pm$ 1.33
DWA 3.99 $\pm$ 0.80 3.38 $\pm$ 0.64 3.57 $\pm$ 0.62 3.59 $\pm$ 0.83
VLM-Social-Nav 4.31 $\pm$ 0.72 4.28 $\pm$ 0.56 4.35 $\pm$ 0.70 4.04 $\pm$ 0.74

### v3

@@ I Introduction
Mobile robots integrated into diverse indoor and outdoor human-centric environments are becoming increasingly prevalent. These robots serve various functions, ranging from package and food delivery [starship] to service [dilligent] and home assistance [astro]. Overall, these roles necessitate interaction with humans and navigating seamlessly through public spaces with pedestrians. In such dynamic scenarios, it is important for the robots to engage in socially compliant interactions and navigation [mirsky2024conflict, mavrogiannis2023core].
@@
Inferring contextually appropriate navigation behaviors is challenging. Humans have various behaviors and the environmental or task contexts cannot be easily categorized [mavrogiannis2023core]. A common strategy to handle the challenge is by learning-based approaches to learn the complicated contexts empirically. Imitation Learning (IL) is a recent emerging paradigm for desired navigation behavior [hirose2023sacson, raj2024targeted]. This approach enables autonomous robots to navigate socially by learning from human demonstrations. Other learning approaches, such as reinforcement learning have also been used to address this problem [kretzschmar2016socially]. While both methods demonstrate promising results in real-world settings, substantial datasets [thorDataset2019, karnan2022scand, nguyenmusohu] for training and reward engineering are required for their successful application and it is hard to generalize.
@@
Figure 1: The trajectories of VLM-Social-Nav (red), DWA (blue), and BC (yellow) approaches in the frontal encountering scenario (left) and the intersection scenario (right). The resulting trajectories show that VLM-Social-Nav demonstrates more socially compliant behavior because it is instructed by a prompt.
Language models are inherently well-suited for contextual understanding but not well applied in social navigation. Recent Large Language Models (LLMs) and Vision-Language Models (VLMs) demonstrate a deep understanding of contextual information and have the potential to perform chain-of-thought [wei2022chain] and common sense reasoning [gpt4v2023]. Those processes are inherent to social navigation, especially the challenges of contextual appropriateness and politeness, which require understanding the task/environmental context and the behavior of humans. This capability has also been evaluated across diverse domains of robotics, including human-like driving scenarios [wen2024road] and autonomous robot navigation [shah2022lmnav]. However, using language models for social navigation is not well explored, the language models suffer from high latency for real-time navigation, and the issue impedes the smoothness and efficiency of human-robot social interaction.
@@
Main Results: In this paper, we present VLM-Social-Nav, a new approach that uses VLMs to interpret contextual information from robot observation to help autonomous robots improve their navigation abilities in human-centered environments. We leverage a VLM to analyze and reason about the current social interaction and generate an immediate preferred robot action to guide an underlying motion planner. We formalize the concept of social cost and the problem definition of social robot navigation suitable for language descriptions. The social cost is defined as how well a robot’s behavior aligns with socially acceptable norms, i.e., the behavior a human would likely exhibit. Our VLM-based scoring module computes the social cost, which is used for a bottom-level motion planner to output appropriate robot actions. To overcome the limitation of existing VLMs’ latency issue, we utilize a state-of-the-art perception model (i.e., YOLO [YOLO2023]) to detect key entities that are used for social interactions (e.g., humans, gestures, and doors) and query a VLM to generate socially compliant navigation behavior and compute the social cost. We demonstrate VLM-Social-Nav in four different indoor scenarios with human interactions. Unlike previous social navigation approaches, VLM-Social-Nav can better navigate through social scenarios by interpreting the situation based on common sense without any dedicated training on a large dataset. Some of our main results include:
@@
  • We propose VLM-Social-Nav, a novel approach for social robot navigation, by integrating VLMs with optimization-based or scoring-based motion planners and a state-of-the-art perception model for better VLM efficiency.
@@
  • We evaluate VLM-Social-Nav in four different real-world indoor social navigation scenarios along with a user study and compare the results with a Dynamic-Window Approach (DWA) [fox1997dynamic] and Behavior Cloning (BC) [pomerleau1988alvinn] method trained on a state-of-the-art large Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand]. VLM-Social-Nav achieves at least $27.38\%$ improvement in average success rate and $19.05\%$ improvement in average collision rate in four scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
More related works and discussions can be found in the technical report of our manuscript [song2024socially].
@@ II-A Safety Requirement of Social Navigation
For social navigation, safety is a basic requirement for interacting with humans and navigating dynamic scenarios [liang2021vo, liang2021crowd]. DWA [fox1997dynamic] is a well-known collision-free navigation method that calculates collision constraints and selects the best feasible action. Velocity-Obstacle (VO)-based approaches are more efficient and can be used to simulate the actions of crowds [best2014densesense], but they do not take uncertainties into account. PRVO [gopalakrishnan2017prvo] and OFVO [liang2021vo] handle the perception uncertainties, but those approaches require a hard threshold for planning. To deal with this issue, learning-based methods empirically train the policies by demonstrations [sun2021motion] or use reinforcement learning to train the robot in a simulator and implement it in real-world scenarios [liang2021crowd, aradi2020survey]. However, learning-based approaches require a significant amount of data or realistic simulators to learn the task. Social interactions are highly nuanced and context-dependent. Simulating these interactions accurately requires sophisticated models of human behavior and interaction dynamics, which is not trivial.
@@ II-B Contextual Appropriateness of Social Navigation
Researchers have developed various methods to incorporate social awareness into mobile robot navigation. Creating such systems is complex, requiring advanced perception and reasoning to navigate environments shared with humans and robots [mirsky2024conflict]. Defining social navigation varies across cultures and platforms. Assessing social compliance, beyond safety, depends on the scenario and requires contextual consideration. Various methodologies are employed to address this challenge, with a significant focus on enhancing learning methods through reinforcement learning, learning from demonstration (particularly by analyzing examples of human trajectories or robots operated by humans), and the utilization of simulated datasets [7539621, li2018role, nazeri2024vanp]. Additionally, various datasets have been collected for this purpose. SCAND [karnan2022scand] and MuSoHu [nguyenmusohu] are two recent large-scale social human navigation datasets in many natural human-inhabited public spaces for robots to learn similar, human-like, socially compliant navigation behaviors. Although extensive research has explored various machine learning techniques, Vision-Language Models (VLMs) have not yet been applied to the social navigation problem, despite their strong potential for contextual analysis.
@@ II-C Large Foundation Models for Navigation
Recent advancements in Language Foundation Models (LFMs) [bommasani2022opportunities], encompassing VLMs and LLMs, show significant potential in robotic navigation. SayCan [Ahn2022] integrates LLMs for high-level task planning. GPT-Driver [mao2023gptdriver] evaluates the performance of GPT-3.5 in simulation for autonomous driving, framing motion planning as a language modeling problem. L3MVN [yu2023l3mvn] constructs semantic maps of environments and utilizes LLMs to reach long-term goals, while LLaDA [li2024driving] enables autonomous vehicles to adapt to diverse traffic rules across regions. LM-Nav [shah2022lmnav] utilizes GPT-3 and CLIP [radford2021learning] to navigate outdoor environments based on natural language instructions, combining language and visual cues for optimal path planning. Despite their powerful capabilities in contextual understanding and commonsense reasoning, language models have not been extensively investigated for social navigation. Our approach proposes a novel method to navigate robots in a socially compliant manner.
@@ III-A Problem Definition
For social robot navigation, humans are no longer perceived only as dynamic obstacles but also as social entities [mirsky2024conflict]. It necessitates integrating social norms into robot behaviors. We define the social robot navigation problem as a Markov Decision Process (MDP): $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C}\rangle$, where $\mathbf{s}=(x,y,\theta)\in\mathcal{S}$ is a state consisting of a robot pose, $\mathbf{a}=(v,w)\in\mathcal{A}$ is an action consisting of a linear and an angular velocity of the robot, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function characterizing the dynamics of the robot, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a cost function. Given a cost function $\mathcal{C}$, the motion planner finds $(v^{*},w^{*})$ that minimizes the expected cost. The cost function takes the following form:
@@
where $\mathcal{C}_{\textrm{goal}}$ encourages movement toward the goal, $\mathcal{C}_{\textrm{obst}}$ discourages collisions with obstacles, and $\mathcal{C}_{\textrm{social}}$ encourages the robot to follow the social norms. $\alpha$, $\beta$, and $\gamma$ are non-negative weights for each cost term.
@@
where $\mathcal{B}$ is a navigation behavior and $\mathcal{B}_{h}$ is a navigation behavior humans would adopt in accordance with social conventions. Minimizing the deviation between them will encourage the robot to emulate socially acceptable human behaviors. While $\mathcal{B}_{h}$ can be obtained through various methods, including large datasets [thorDataset2019, karnan2022scand, nguyenmusohu], we leverage the power of a VLM to compute appropriate behavior based on its rich contextual understanding and nuanced interpretations from perceived images and given prompts. We elaborate further in Section III-C.
@@ III-B VLM-based Social Navigation Architecture
Fig. 2 highlights the overview of VLM-Social-Nav. Our approach is based on an autonomous navigation system that integrates a perception layer with an optimization-based motion planner. The motion planner processes sensor inputs and generates a robot action that minimizes the cost function $C$.
While LiDAR detects geometric information useful for obstacle avoidance, RGB images provide contextual details of the current environment. They contain rich information crucial for social navigation. To enhance navigation capabilities within social contexts, we propose a VLM-based scoring module. VLMs excel in contextual understanding, interpreting scenes not solely based on visual features but also considering social dynamics[duan2024manipulate]. VLMs generate socially appropriate robot actions based on current observations and input instructions. Our VLM-based scoring module then calculates a cost term to be used by the motion planner.
@@ 5       $e\leftarrow\text{Perception Model}(\mathcal{I})$;
6       forall possible actions $\mathbf{a}$ do
7             $\mathcal{C}_{\textrm{social}}\leftarrow 0$;
@@ 9                   $\mathcal{B}_{h}=\text{VLM}(\mathcal{I},\mathcal{P},\mathbf{a})$;
10                   $\mathcal{C}_{\textrm{social}}=\text{VLM-based scoring}(\mathcal{B}_{h})$;
11
12             Calculate the total coast $\mathcal{C}=\alpha\cdot\mathcal{C}_{\textrm{goal}}+\beta\cdot\mathcal{C}_{\textrm{obst}}+\gamma\cdot\mathcal{C}_{\textrm{social}}$;
13
@@ 16       Update robot position $\mathbf{p}_{r}$;
Algorithm 1 VLM-based Social Navigation
III-C VLM-based Scoring Module
VLM plays a crucial role in VLM-Social-Nav in inferring immediate socially compatible navigation behavior $\mathcal{B}_{h}^{t+1}$ based on its pre-trained large internet-scale dataset:
where $\mathcal{I}^{t}$ is an RGB image from the robot view at time $t$, $\mathcal{P}$ is a textual prompt, and $\mathbf{a}^{t}$ is a current robot action at time $t$. Inspired by In-Context Learning (ICL), our prompt $\mathcal{P}$ is designed to leverage the VLM’s reasoning abilities through zero-shot examples. This approach offers an interpretable interface, mirroring human reasoning and decision-making processes, without extensive training [min2022rethinking].
where $w_{l}$ and $w_{a}$ are non-negative weights. Given all the cost terms, our low-level optimization-based motion planner finds the robot action $(v^{*},w^{*})$ that minimizes the cost.
Fig. 3 shows an example prompt $\mathcal{P}$ used in our experiment. We provide a high-level task description along with an image $\mathcal{I}^{t}$ captured from the robot’s perspective. Furthermore, the current robot action $\mathbf{a}^{t}=(v^{t},w^{t})\in\mathcal{A}$ is provided. The angular velocity is mapped into corresponding directional instructions based on predefined categories (i.e., positive values correspond to left, values near zero to straight, and negative values to right). Supplementary instructions regarding walking etiquette are included. Although the VLM demonstrates proficient navigation abilities in the absence of explicit instructions, offering reasoning guidelines enhances its decision-making processes [min2022rethinking]. These guidelines not only facilitate comprehensive reasoning and judgment within the VLM but also enable the robot to adapt to specific rules more effectively. For example, in a country where it’s customary to walk on the left, we can rephrase the prompt as “Move to the left when passing by another person.”
@@ III-C VLM-based Scoring Module
Figure 3: An example input image ($\mathcal{I}^{t}$) and prompt ($\mathcal{P}$) used in VLM-Social-Nav. Parameterized inputs ($\mathbf{a}^{t}$) are highlighted in blue. Formatted outputs specifying the heading ($\delta_{d}$) and the speed ($\delta_{s}$) are highlighted in red. The example input data is one of the frontal approach scenarios from MuSoHu [nguyenmusohu].
@@
Figure 4: Qualitative Results: the robot navigation behaviors with VLM-Social-Nav for four social navigation scenarios: (a) Frontal Approach, (b) Frontal Approach with Gesture, (c) Intersection, and (d) Narrow Doorway. The solid gray arrow shows the participant’s path. The solid red arrow shows the robot’s path. The red and gray dashed arrows show the robot’s and participant’s paths respectively, after a stop motion. A caption on the top left shows the result from the VLM.